### v1

# Reinforcement learning to learn quantum states for Heisenberg scaling accuracy

Abstract
Learning quantum states is a crucial task for realizing the potential of quantum information technology. Recently, neural approaches have emerged as promising methods for learning quantum states. We propose a meta-learning model that employs reinforcement learning (RL) to optimize the process of learning quantum states. For learning quantum states, our scheme trains a Hardware efficient ansatz with a blackbox optimization algorithm, called evolution strategy (ES). To enhance the efficiency of ES, a RL agent dynamically adjusts the hyperparameters of ES. To facilitate the RL training, we introduce an action repetition strategy inspired by curriculum learning. The RL agent significantly improves the sample efficiency of learning random quantum states, and achieves infidelity scaling close to the Heisenberg limit. We showcase that the RL agent trained using $3$-qubit states can be generalized to learning up to $5$-qubit states. These results highlight the utility of RL-driven meta-learning to enhance the efficiency and generalizability of learning quantum states. Our approach can be applicable to improve quantum control, quantum optimization, and quantum machine learning.
I introduction
Learning quantum states has emerged as a significant problem in the field of quantum computing and machine learning [1, 2]. As a tool for analyzing and benchmarking quantum systems, learning quantum states is essential for realizing quantum information technology [3]. The purpose of learning quantum states is to estimate quantum states with high fidelity based on samples obtained by measuring copies of states. Quantum state tomography (QST) is the standard method for learning quantum states [4, 5], but this method becomes impractical for large systems as the number of measurements required scales exponentially with the size of system. This limitation fuels research directions to devise various strategies for QST using matrix product state [6, 7], permutation invariance of state [8], shadow tomography [9], and Bayesian estimation [10].
With the advent of deep learning, there has been a growing body of work using a neural network-based model for learning quantum states [11, 12, 13, 14, 15, 16, 17, 18, 19]. Generative models provide sample-efficient methods for learning many-body quantum states based on restricted Boltzmann machine [20], variational autoencoder [21], and recurrent neural network [22]. These models are trained with unsupervised learning. Using supervised learning, a model of feed-forward neural network enhances QST by removing errors [23]. The power of these approaches stems from the expressivity of neural network, the ability to represent intricate features of quantum states in terms of a small number of parameters [24]. In the framework of quantum machine learning, a model based on a quantum neural network has also been employed for learning quantum systems [25, 26, 27, 28, 29]. The quantum model can be implemented with parameterized quantum gates [30, 31], making it well suited for learning states of quantum computers. Moreover, it is known that a quantum neural network can have higher expressive power than classical neural networks of comparable size [32, 33].
Meanwhile, the generalizability, the ability to generalize to unexperienced tasks, is one of the hallmarks of machine learning model. However, some of the models for learning quantum states are difficult to generalize as each model is tailored to learning a specific quantum state. Learning models with the generalizability can be found in attempts to improve a quantum optimization algorithm using a deep learning technique [34, 35, 36, 37]. These works employ a hybrid approach that leverages a classical model to learn a quantum model. Specifically, Verdon et al. propose a model based on a recurrent neural network to initialize the parameters of quantum approximate optimization algorithm (QAOA) circuit [35]. Similar work by Wilson et al. uses a recurrent neural network model to learn the gradient of objective function of QAOA [36]. Khairy et al. show that a feed-forward neural network model can be used to learn the hyperparameters of QAOA with reinforcement learning [37]. They show that the hybrid models can be generalized to solving problems of higher dimension than the dimension of problems for training. The deep learning technique in the previous works can be categorized as meta-learning, also called learning to learn, which is employed to enable a model to adapt to unexperienced tasks [38, 39]. Meta-learning methods for learning quantum states have also been proposed, where a classical model is used to learn another classical model [40] or Bayesian estimation algorithm [41]. For the task of learning quantum states, the validity and generalizability of meta-learning models in the hybrid setting remains an area of exploration.
In this work, we propose a meta-learning approach which leverages reinforcement learning (RL) to learn quantum states. The meta-learning model is based on the quantum and classical neural networks. We consider a Hardware efficient ansatz as the quantum neural network model [42]. The quantum model learns a quantum state with a blackbox optimization algorithm, called evolution strategy (ES) [43]. To enhance the process of learning quantum states, a RL agent consisting of classical neural networks dynamically adjusts the hyperparameters of ES. We use single-shot measurements to train the quantum and classical models. It has been shown that using single-shot measurement can exhibit high shot efficiency in learning quantum states [25, 26]. To facilitate the RL, we introduce action repetition strategy inspired by curriculum learning, a technique for training a learning model by gradually increasing difficulty of problems [44]. The trained RL agent significantly improves the shot efficiency of learning quantum states, and enables the infidelity $\bar{f}$ to approach the scaling of Heisenberg limit for the total number of measurement shots ${\cal C}$ [45], i.e., $\bar{f}\sim{\cal O}({\cal C}^{-1})$. We generalize the RL agent to learning quantum states of higher dimension than the dimension of quantum states used for training.
The remainder of this paper is organized as follows. We explain the scheme of meta-learning in Section II. In section III, we present results of meta-learning for random pure states, and introduce the action repetition strategy. We also investigate the generalizability of our model. We give discussion and conclusion in Section IV.
Figure 1: Schematic of reinforcement learning to learn quantum states. (a) In the environment, the evolution strategy trains a Hardware efficient ansatz (HEA) in (b) to learn an input state $\left|\psi\right>$. The observation $o_{t}$ is the success count $\cal C$, and the HEA is trained to increase the success count. The RL agent learns the action $a_{t}$, the hyperparameters of evolution strategy, to enhance the learning quantum state. To enable the learning quantum states to be finished quickly, we give a penalty $r_{t}=-1$ for each time step $t$, and no penalty $r_{t}=0$ for the time step when the learning quantum states is finished. The rollout datasets are generated by a policy-guided tree search, and stored in the data buffer. We use the Actor-Critic algorithm to train the RL agent with a subdataset $\cal B$ randomly drawn from the buffer. (b) HEA circuit. $U_{3}$ is a generic single-qubit gate $\exp({i\vec{\theta}\cdot\vec{\sigma}/2})$, where $\vec{\theta}=(\theta_{x},\theta_{y},\theta_{z})$ and $\vec{\sigma}$ is a vector of Pauli operators $(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$. For 1-qubit states, we use a single $U_{3}$ gate. For $N$-qubit states, we use a HEA of $L$ layers which has $3N(L+1)$ parameters. (c) Feed-forward neural networks for the RL agent. The input of each network is the observation $o_{t}$. $Q_{\phi_{1}}$ is the observation-value function and $\pi_{\phi_{2}}$ is the policy of action.
II Meta-learning: RL to learn
The schematic of meta-learning model is illustrated in Fig. 1 (a). The environment contains an evolution strategy and a quantum system. As shown in Fig. 1 (b), the quantum system consists of a quantum state $\left|\psi\right>$, a Hardware efficient ansatz (HEA) $\hat{U}(\vec{\theta})$, and a measurement. The evolution strategy trains the HEA to learn an input quantum state based on the measurement. The RL agent learns the hyperparameters of evolution strategy using two classical models in Fig. 1 (c), called actor and critic. We describe our RL scheme with a Markov decision process.
Markov decision process (MDP) is a framework to describe a sequential interaction that takes place between an environment and an agent [46]. The agent takes an action $a$ based on observation $o$ about the environment. As a result of an action, the state of environment is changed, and the agent receives a reward $r$ from the environment. A purpose of agent is to learn the policy of action $\pi$ to maximize cumulative rewards (see Appendix A for a detailed description).
In determining approach of agent for formulating and solving a MDP, observability of the environment is a significant factor to consider. When the elements of MDP are fully known, that is, assuming full observability of environment, an agent can derive the exact policy based on Bellman’s principle of optimality [47]. However, in many scenarios, an agent cannot completely specify the state of environment and its dynamics due to limited (or partial) observability. Partially observable MDP (POMDP) is a generalized framework accommodating such limited observability [48]. To solve a POMDP, an agent can take a heuristic method through an iterative process. A deep RL is known as a method for learning the heuristic policies in the absence of full observability [49].
The problem we consider is the case that the RL agent does not have the full observability of environment. A quantum state transformed by the HEA, $\hat{U}(\vec{\theta})\left|\psi\right>$, represents the state of environment. However, the exact state of environment is unknown to the agent, and the single-shot measurement can extract only a partial information about the quantum state. Thus, we can formulate our problem with the description of POMDP [48, 50, 51] defined by a tuple $\left({\cal S},{\cal A},{\cal T},{\cal R},\Omega,{\cal O}\right)$. The elements of POMDP are explained in Appendix B.
The major challenge in applying RL to quantum systems is the partial observability by quantum measurements [50]. To apply RL for control of quantum systems, there are attempts to utilize additional information about quantum states such as fidelity [52, 53, 54, 55, 56, 57], but computing this quantity is typically expensive, comparable to learning quantum states by the standard QST methods. In contrast, the RL method proposed by Sivak et al. learns the control policy based on observable quantities [51]. In the same spirit, we train our RL agent using the observation $o$ obtained by the measurement. In the following subsections, we describe in detail the configuration of environment, agent, and RL algorithm.
II.1 Environment
Learning a quantum state.—Consider an unknown $N$-qubit state $\left|\psi\right>$ as the input of HEA $\hat{U}(\vec{\theta})$. We measure the quantum state transformed by the HEA, $\hat{U}(\vec{\theta})\left|\psi\right>$, with a measurement $M=\{\hat{M}_{s},\hat{M}_{f}\}$ prepared in known bases, where
$\hat{M}_{s}=|s\rangle\langle s|,\quad\hat{M}_{f}=I-\hat{M}_{s}.$ (1)
The outcome $s$ stands for success and $f$ does fail. We perform the measurement until a fail outcome appears, and count the number of success outcomes before the fail outcome. If a fail outcome appears, we change the parameters of quantum circuit based on the number of success outcomes. Otherwise, we retain the parameters. This measurement scheme is the same as that of single-shot measurement learning (SSML) [25, 26]. While SSML uses a random search algorithm to update the parameters, we use the evolution strategy which will be explained later.
We refer to the number of consecutive success outcomes as success count. The success count is determined by the probability of success outcome $p_{s}=\langle\psi|\hat{U}^{\dagger}(\vec{\theta})\hat{M}_{s}\hat{U}(\vec{\theta})|\psi\rangle$. Thus, the chance to obtain ${\cal C}$ success count is governed by a geometric distribution
$p({\cal C})=p_{s}^{\cal C}(1-p_{s}).$ (2)
The average value of success count is then given by $\langle{\cal C}\rangle=\sum_{{\cal C}=0}^{\infty}{\cal C}p({\cal C})=p_{s}/(1-p_{s})$. If $p_{s}\rightarrow 1$, $\langle{\cal C}\rangle\rightarrow\infty$, and, if $p_{s}=0$, $\langle{\cal C}\rangle=0$. The success count is the observation $o$, and the geometric distribution determines the transition probability of observation.
We train the HEA until the success count achieves a high value so that the HEA transforms the input quantum state close to the basis of success outcome. We set a target success count ${\cal C}_{\text{target}}$, and have the training process stop if a success count reaches or exceeds the target value, i.e., ${\cal C}\geq{\cal C}_{\text{target}}$. The value of target success is related to the accuracy of reconstructed quantum state.
Quantum state reconstruction.—After the training, we can reconstruct the input quantum state through the trained HEA as
$\left|\psi_{\text{est}}\right>=\hat{U}^{\dagger}\left(\vec{\theta}_{\text{train}}\right)\left|s\right>.$ (3)
To assess the accuracy of the reconstructed state $\left|\psi_{\text{est}}\right>$, we choose the infidelity as a figure of merit [45] defined by
$\bar{f}:=1-|\langle\psi|\psi_{\text{est}}\rangle|^{2}$ (4)
The infidelity depends on the target success count ${\cal C}_{\text{target}}$ we initially set; The larger the target success count, the better the accuracy of the estimation.
Optimizer.—To train the HEA, we employ an evolution strategy (ES) [58]. An ES consists of following steps: (i) assuming a probability distribution for sampling, (ii) sampling parameters around a current parameter, (iii) evaluating the parameter samples, and (iv) updating the current parameter based on the evaluations. In our problem, the ES uses a multivariate Gaussian distribution to sample parameters [43], and the evaluation corresponds to measuring the success count.
The purpose of ES is to find the parameters of HEA $\vec{\theta}$ to achieve the halting condition ${\cal C}(\vec{\theta})\geq{\cal C}_{\text{target}}$. For this purpose, we consider the expectation of success count as an objective function of ES
$J(\vec{\theta}):=\frac{1}{{\cal C}_{\text{target}}}\mathbf{E}_{\vec{\theta}\sim p(\vec{\theta})}\left[{\cal C}(\vec{\theta})\right],$ (5)
where the target success count ${\cal C}_{\text{target}}$ is introduced in the denominator to normalize the objective function. $p(\vec{\theta})$ is the multivariate Gaussian distribution ${\cal N}(\vec{\theta},\sigma^{2}I)$ of mean $\vec{\theta}$ and covariance matrix $\sigma^{2}I$, where $\sigma\in(0,\infty)$ is a hyperparameter to determine the range of sampling over the space of parameters, and $I$ is the identity matrix.
To increase the expected success count, the parameters are updated by following rule
$\vec{\theta}\leftarrow\vec{\theta}+\eta\nabla_{\vec{\theta}}\tilde{J}(\vec{\theta}),$ (6)
where $\eta\in[0,\infty)$ is the learning rate, and $\nabla_{\vec{\theta}}\tilde{J}(\vec{\theta})$ is an estimator of gradient. Applying the reparameterization trick [43], the ES estimates the gradient by drawing $\{\vec{\epsilon_{i}}\}_{i=1}^{k}$ from a distribution ${\cal N}(\vec{0},I)$, and evaluating $k$ parameter samples $\{\vec{\theta}+\sigma\vec{\epsilon}_{i}\}_{i=1}^{k}$. The estimator of gradient is given by
$\nabla_{\vec{\theta}}\tilde{J}(\vec{\theta})=\frac{1}{{\cal C}_{\text{target}}\sigma k}\sum_{i=1}^{k}{\cal C}(\vec{\theta}+\sigma\vec{\epsilon}_{i})\vec{\epsilon}_{i}.$ (7)
{algorithm}
[t] Evolution strategy with RL agent
1:Set an input quantum state $\left|\psi_{1}\right>$, the HEA $\hat{U}(\vec{\theta}_{0})$, the target success count ${\cal C}_{\text{target}}$, the number of samples $k$, and the RL agent $\pi$
2:for $t=1,2\cdots,t_{\text{max}}$ do
3:     Measure ${\cal C}(\vec{\theta}_{t})$ and $o_{t}\leftarrow{\cal C}(\vec{\theta}_{t})$
4:     Draw $\sigma_{t}$ and $\eta_{t}$ from the agent, i.e., $(\sigma_{t},\eta_{t})\sim\pi(a_{t}|o_{t})$
5:     Sample $\vec{\epsilon}_{1},\cdots,\vec{\epsilon}_{k}\sim{\cal N}(0,I)$
6:     Measure ${\cal C}\left(\vec{\theta}_{t}+\sigma_{t}\vec{\epsilon}_{i}\right)$ for $i=1,\cdots,k$
7:     if any ${\cal C}\geq{\cal C}_{\text{target}}$ then
8:         break
9:     end if
10:     $\vec{\theta}_{t+1}\leftarrow\vec{\theta}_{t}+\eta_{t}\nabla_{\vec{\theta}}\tilde{J}(\vec{\theta}_{t})$triangleright Eq. (7)
11:     $\left|\psi_{t+1}\right>\leftarrow\hat{U}(\vec{\theta}_{t+1})\left|\psi_{1}\right>$
12:end for
13:${\cal C}_{\text{total}}\leftarrow$ the sum of success counts
14:$t_{H}\leftarrow$ $t$
15:$\vec{\theta}_{\text{train}}\leftarrow\vec{\theta}$ s.t. ${\cal C}(\vec{\theta})\geq{\cal C}_{\text{target}}$
The hyperparameter of sampling range, also called step size, is a significant factor to determine the performance of ES. The sampling range decides the range of search: A larger sampling range induces more exploration on the parameter space. A smaller sampling range focuses on exploitation, searching in a narrower region to enhance solution. In advanced variants of ES such as covariance matrix adaptation ES [59], the sampling range and learning rate are dynamically adjusted, and their combination determines the covariance matrix (or the inverse of Hessian matrix) which reflects the local geometry of objective function. In our scheme, the sampling range $\sigma$ and learning rate $\eta$ are dynamically changed according to the policy $\pi$.
The pseudocode of ES is shown in Algorithm II.1. Unlike the two hyperparameters $\sigma$ and $\eta$, we fix the number of samples $k$ to a single value to keep the statistical uncertainty of the gradient estimator at a certain degree during the execution of ES. ${\cal C}_{\text{total}}$ is the total success count used, and the halting time, $t_{H}$, is the time steps taken until the training is finished. Note that the time here means that how many loops the ES takes until the success count reaches the halting condition ${\cal C}\geq{\cal C}_{\text{target}}$. In a real experiment, the total success count is a quantity which is proportional to physical runtime for the training.
An optimizer typically used to train a quantum neural network is a gradient method [60, 61], but it can suffer the gradient vanishing problem, called barren plateau [62, 63], making the training intractable. While the gradient vanishing effect cannot be fully alleviated [64], recent studies have shown that an ES can be useful for training some cases of quantum neural networks [65, 66]. The gradient methods typically require more evaluation points than parameters to obtain directional derivatives [60]. In contrast, we will show that the ES can train the HEA with the fewer evaluations than the parameters of HEA.
Figure 2: Results of training RL agent for the target success count ${\cal C}_{\text{target}}=10^{4}$. For each RL training episode $T$, we perform learning a thousand random states. $\langle{\cal C}_{\text{total}}\rangle$ is the average of total success count. $\langle\bar{f}\rangle$ is the average infidelity of learned states. $\langle t_{H}\rangle$ is the average halting time steps of evolution strategy. (a) The red, green, and blue solid lines are the training curves for learning $1$-, $2$-, and $3$-qubit states with the $16$ actions. The black solid line is obtained by using the extended action space for learning $2$-qubit states. The dashed lines are baseline values. (b) The policy of trained RL agent.
II.2 RL agent
The RL agent consists of an actor and a critic, both implemented by feed-forward neural networks [see Fig. 1 (c)]. The both networks use three hidden layers. For each hidden layer, the critic has hundred nodes, and the actor has fifty nodes. We use ReLu for activation functions. The actor is to learn a policy of action. We construct the action space ${\cal A}$ by selecting $m$ values from each of the hyperparameter spaces ${\cal A}_{\sigma}=[0,\infty)$ and ${\cal A}_{\eta}=[0,\infty)$. So the number of actions the agent can choose from is $m^{2}=\left|\cal A\right|$. The output of actor is the policy $\pi_{\phi_{2}}=p(\sigma,\eta|o,\phi_{2})$, the probability distribution defined over the action space ${\cal A}$. The critic outputs an estimate of observation-value function $Q_{\phi_{1}}(o)$.
We train the agent to recommend the hyperparmaeters $\sigma$ and $\eta$ that enables the learning quantum states to be finished quickly. This is accomplished by giving a penalty of $-1$ for each time step, and no penalty for the time step when the learning quantum states is finished. In other words, the reward is given by
$r_{t}=\begin{cases}~{}~{}0&\text{if}~{}~{}t=t_{H}\\ -1&\text{otherwise}.\end{cases}$ (8)
The empirical cumulative reward is then given by
${\cal R}_{t}=\sum_{t}^{{t_{H}}}r_{t}={t}-{t_{H}}.$ (9)
For the training, we use the Actor-Critic algorithm, and collect training datasets using a tree search guided by the policy. See Appendix C for detailed description.
Table 1: The settings used to train the RL agent. The space of action is given by ${\cal A}={\cal A}^{4}_{\sigma}\times{\cal A}^{4}_{\eta}$. The action repetition time $t_{\text{rep}}$ is determined by the parameters $(t_{l},t_{u},T_{\text{th}})$. For $4$- and $5$-qubit, we perform the learning random pure states with the RL agent trained using the $3$-qubit states.
No. qubits $N$ Quantum neural network No. samples $k$ The space of sampling range ${\cal A}_{\sigma}^{4}$ The space of learning rate ${\cal A}_{\eta}^{4}$ Action repetition strategy $(t_{l},t_{u},T_{\text{th}})$
$1$ $U_{3}$ 5 $\{1.0,0.1,0.01,0.001\}$$\{1.0,0.1,0.01,0.001\}$ $(1,50,100)$
$2$ HEA, $L=1$ 10 $\{1.0,0.1,0.01,0.001\}$$\{1.0,0.1,0.01,0.001\}$ $(80,800,100)$
$3$ HEA, $L=5$ 30 $\{0.1,0.01,0.001,0.0001\}$$\{1.0,0.33,0.01,0.033\}$ $(300,2000,500)$ [^1] We obtain similiar results for $t_{l}=200$.
$4$ HEA, $L=10$ 100 $\{0.1,0.01,0.001,0.0001\}$$\{1.0,0.33,0.01,0.033\}$ $t_{\text{rep}}=300$
$5$ HEA, $L=10$ 100 $\{0.1,0.01,0.001,0.0001\}$$\{1.0,0.33,0.01,0.033\}$ $t_{\text{rep}}=300$
III Results
III.1 Training RL agent
[…]
We present the setting of HEA, the number of samples, and the space of action in Table 1. While the number of parameters of unitary operator required to prepare an $N$-qubit state is $4^{N}$, we show that the HEA can learn the random quantum states with fewer parameters: For the learning $2$- and $3$-qubit states, we use the HEAs which have $12$ and $54$ parameters, respectively. We will also show that HEA of $10$ layers can learn random $4$- and $5$-qubit states. Moreover, to estimate the gradient of HEAs, we use fewer samples than the parameters of HEA circuits. For $2$- and $3$-qubit states, the number of samples is set to $10$ and $30$, respectively.
Figure 3: (a) The scheme of action repetition strategy (ARS). If the ARS is not applied, the agent $\pi$ draws a new action $a_{i}$ based on the observation $o_{t}$ at each time step. With the ARS, the agent repeats an action chosen during $t_{\text{rep}}$ time steps. The example shows the case of $t_{\text{rep}}=3$. (b) Learning random $1$-qubit states with and without ARS. The red solid line is obtained by using the ARS with setting $(t_{u},t_{l},T_{\text{th}})=(50,1,100)$. The gray solid line is obtained without the ARS. The dashed line is the baseline value (ES only).
The results of training RL agent are illustrated in Fig. 2 (a). For these results, we set the target success count to ${\cal C}_{\text{target}}=10^{4}$, and randomly initialize the parameters of HEAs. We compare these results to the performance obtained without the RL, i.e., using only a single action for the entire process of learning quantum states. We regard results obtained from an action which gives the lowest total success count as baseline values, and train the RL agent until it outperforms the baseline values. The baselines are selected based on data in Appendix. D.
At the end of RL training ($T\approx 1500$), the average of total success counts required to learn the $1$-, $2$-, and $3$-qubit states are $3.123\times 10^{4}$, $3.767\times 10^{5}$, and $2.363\times 10^{6}$, which are about $14\%$, $28\%$, and $34\%$ lower than the baseline values, respectively. Using the RL, we obtain the (in)fidelity similar to the baselines, except for the case of $3$-qubit. The average fidelity is given by $0.99994$, $0.99972$, and $0.99946$ for $1$-, $2$-, and $3$-qubit states, respectively. In the early episodes of RL, the average halting time $t_{H}$ of learning $1$-, $2$-, and $3$-qubit states is about $10^{3}$, $10^{4}$, and $1.5\times 10^{4}$, and these are reduced to $19$, $660$, and $3409$ at the end of episode, respectively. For these results, we use the $16$ actions defined in Table 1.
For the $2$-qubit states, we also consider an extended action space defined by ${\cal A}:={\cal A}^{13}_{\sigma}\times{\cal A}^{13}_{\eta}$, where
${\cal A}^{13}_{\sigma}$ $=$ $\left\{\sigma_{i}~{}\bigg{|}~{}\sigma_{i}=10^{-0.25(i-1)}~{}\text{for}~{}i=1,\ldots,13\right\}$ (10)
${\cal A}^{13}_{\eta}$ $=$ $\left\{\eta_{i}~{}\bigg{|}~{}\eta_{i}=10^{-0.25(i-1)}~{}\text{for}~{}i=1,\ldots,13\right\}.$
The number of actions in the extended space is $169$. Except for the number of actions, we use the same settings for the $2$-qubit case listed in Table. 1. The result is shown in the rightmost column of Fig. 2 (a). By extending the action space, we can attain the lower values of average total success count and infidelity than the case of using the $16$ actions. The average total success count is $3.593\times 10^{5}$ and the average fidelity is $0.9998$. To achieve these results, on average, $433$ time steps are required. While the overall performance of learning $2$-qubit states is improved by extending the action space, the agent requires more episodes to be trained.
Fig. 2 (b) shows the policy of trained actor. These policies imply that the trained RL agent takes a strategy; if the success count is small, the agent recommends the large values for $\sigma$ and $\eta$ to enable the ES to search in a wide range of parameter space; if the success count is large, the agent reduces the values of recommended hyperparameter to make the learning quantum states to converge.
Figure 4: Learning random pure states with trained RL agent. We consider the target success counts from $10^{1}$ to $10^{4}$. The dots in (a)–(c) and (e) represent the average total success count $\langle{\cal C}_{\text{total}}\rangle$ and the average infidelity $\langle\bar{f}\rangle$ for the learning $1$-qubit (red), $2$-qubit (green), and $3$-qubit (blue) states. We denote the results obtained using the trained RL agent, referred to as ES$+$RL. The black triangles are the average values obtained without the RL agent, which are denoted by ES. For the results of ES, we use an action $a_{\text{base}}$ which gives the baseline values. SSML denotes the results obtained by the single-shot measurement learning. (d) $\Delta\langle{\cal C}_{\text{total}}\rangle=\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES}}}-\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES+RL}}}$ is the amount of average total success counts reduced by using the trained RL agent. (e) The purple and black dots are results of generalizing the RL agent trained using $3$-qubit states to learning random $4$- and $5$-qubit states, respectively. The dashed lines represent the scaling of statistical limit, $\bar{f}\sim{\cal O}({\cal C}^{-1}_{\text{total}})$.
III.2 Action repetition strategy
To obtain these results, we use a method, called action repetition strategy (ARS), that repeats an action given by the RL agent for a certain period of time steps $t_{\text{rep}}$. The schematic of ARS is illustrated in Fig. 3 (a). For each RL episode $T$, $t_{\text{rep}}$ is given by
$t_{\text{rep}}$ $=$ $\max\left(\left\lceil{t_{u}-\frac{T}{T_{\text{th}}}\left({t_{u}-t_{l}}\right)}\right\rceil,t_{l}\right),$ (11)
where $t_{u}$ and $t_{l}$ are the upper and lower value of $t_{\text{rep}}$, respectively, and $t_{u}>t_{l}$. When $T=0$, the action repetition time is $t_{u}$, and it decreases until the RL episode $T$ reaches $T_{\text{th}}$. After $T_{\text{th}}$, $t_{\text{rep}}$ saturates to $t_{l}$.
The ARS is a method inspired by curriculum learning, a technique that trains a machine learning model by gradually increasing the difficulty of problems [44]. Note that the curriculum learning belongs to the category of transfer learning, which reuses a model or knowledge obtained from a task to learn a related task [70, 71, 72]. One of factors that determines the difficulty of RL problem is the depth of decision process. In general, the deeper the decision process, the more difficult the problem is considered to be. The depth of decision process of our problem $t_{H}$ in the early episodes is about $10^{3}$, $10^{4}$, and $1.5\times 10^{4}$ for the learning $1$-, $2$-, and $3$-qubit states, respectively. With the ARS, we can reduce the depth of decision process to $\left\lfloor t_{H}/t_{\text{rep}}\right\rfloor$, which is shallower than the original depth. Specifically, according to the settings in Table 1, the decision process can be reduced to tens of steps in the beginning of the RL training. This leads to the significant decrease of difficulty of problem. As the RL training proceeds, the difficulty of problem increases again with the decrease of $t_{\text{rep}}$. Thus, by controlling the action repetition time, the ARS enables the RL agent to experience problems of various difficulty levels during the training process.
For the random $1$-qubit states, we compare the performance obtained with and without the ARS in Fig 3 (b). With the ARS, the RL agents can converge faster than the case without the ARS. For the random $2$- and $3$-qubit states, the training RL agent without the ARS requires computational resources far exceed our capacity. In some intervals of RL episodes, the difficulty of our problem does not gradually increase since the halting time step $t_{H}$ also decreases together with the action repetition time. More sophisticated curriculum for the repetition time may improve the performance of meta-learning, but it is beyond the scope of this work and remains an open question.
Figure 5: Learning an entangled $5$-qubit pure state $\hat{\varrho}_{SC}$ with the RL agent trained using the random $3$-qubit states. (a) Re$\left(\hat{\varrho}_{SC}\right)$ represents the real parts of density matrix of $5$-qubit state. The imaginary parts of density matrix are zero. (b) The reconstructed state $\hat{\varrho}_{\text{est}}$ of the $5$-qubit state. Re$\left(\hat{\varrho}_{\text{est}}\right)$ and Im$\left(\hat{\varrho}_{\text{est}}\right)$ represent real and imaginary parts of density matrix $\hat{\varrho}_{\text{est}}$, respectively. The reconstructed state achieves the fidelity $f\approx 0.9989$. (c) The changes of infidelity during the entire learning process.
III.3 The scaling of infidelity
Using the RL agents trained with the target success count ${\cal C}_{\text{target}}=10^{4}$, we perform the learning a hundred random pure states by varying the target success counts from $10^{1}$ to $10^{4}$. The results are shown in Fig. 4 (a)–(c). We compare these results to the baseline values obtained by an action $a_{\text{base}}$ (see Appendix D). Fig. 4 (d) shows the amount of total success counts saved by the RL agent according to the increase of the number of qubits $N$ and target success count ${\cal C}_{\text{target}}$. The amount of total success counts saved is represented by $\Delta\langle{\cal C}_{\text{total}}\rangle:=\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES}}}-\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES+RL}}}$, where $\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES}}}$ is obtained using $a_{\text{base}}$ without the RL agent, and $\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES+RL}}}$ is the average total success count obtained with the trained RL agent. In the target success count $10^{4}$, $\Delta\langle{\cal C}_{\text{total}}\rangle$ is given by $3.274\times 10^{4}$, $1.995\times 10^{5}$, and $2.818\times 10^{6}$ for $1$-, $2$-, and $3$-qubit states, respectively. This leads to enhancement in the scaling of average infidelity $\bar{f}$ versus the total success counts ${\cal C}_{\text{total}}$. The scaling nearly achieves the Heisenberg limit [45], $\bar{f}\sim{\cal O}({\cal C}^{-1}_{\text{total}})$, which the standard QST method cannot attain without additional information about quantum states [73].
We curve fit the results of Fig. 4 (a)–(c) with $\bar{f}=\alpha{\cal C}_{\text{total}}^{-\beta}$. The scaling factor is given by $\beta\approx 0.948$, $1.161$, and $1.184$ for the learning $1$-, $2$-, and $3$-qubit states, respectively. Our method shows the similar results to previous works on learning $1$-qubit states [74, 75, 76, 25, 26], where the scaling factors lie in the range $\beta\in[0.9,1.0)$. Typically, the SSML shows the scaling of $\beta\approx 0.972$ in our simulation, which is consistent with the results in their works [25, 26]. (In the multi-qubit cases, we fail to find hyeperameter settings of SSML to achieve the optimal scaling.) Our result obtained with the RL agent achieves the similar scaling factor while using fewer total success count than the SSML as shown in Fig. 4 (a).
III.4 Generalization of RL agent
We apply the RL agent trained with the random $3$-qubit states to learning a hundred random $4$- and $5$-qubit states. The results are illustrated in Fig. 4 (e). For the target success count ${\cal C}_{\text{target}}=10^{4}$, the average of total success count and infidelity are given by $\langle{\cal C}_{\text{total}}\rangle\approx 7.261\times 10^{6}$ and $\bar{f}\approx 6.561\times 10^{-4}$ for $4$-qubits, and $\langle{\cal C}_{\text{total}}\rangle\approx 1.291\times 10^{8}$ and $\bar{f}\approx 9.311\times 10^{-4}$ for $5$-qubits. The scaling factor of infidelity achieves $\beta=-1.189$ and $\beta=-0.829$ for $4$- and $5$-qubit, respectively. For the multi-qubit cases including $2$- and $3$-qubit results, our method gives the similar results to the self-guided quantum state tomography (SGQT) which shows the scaling factors in the range $\beta\in(0.80,1.05)$ up to $10$-qubits [76].
By using the trained RL agent, we learn a specific $5$-qubit state, and reconstruct the learned state. As illustrated in Fig. 5 (a), the density matrix of $5$-qubit state $\hat{\varrho}_{SC}$ is given by a Shen-Castan matrix of which element is determined by a real-valued function $\exp[-(\left|m-16.5\right|+\left|n-16.5\right|+1)/10]/K$ for $m,n=1,\ldots,32$ and $K=0.00366$. (A Shen-Castan matrix is a filer used for edge detection in image processing [77].) This quantum state is pure, i.e., $\mbox{Tr}\hat{\varrho}_{SC}^{2}=1$, and the five qubits are entangled. For a reduced state of $i$-th subsystem $\hat{\varrho}_{i}$, Von Neumann entropy ${S}_{i}:=-\mbox{Tr}\hat{\varrho}_{i}\log_{2}\hat{\varrho}_{i}$ has nonzero value: ${S}_{1}=0.640$, ${S}_{2}=0.547$, ${S}_{3}=0.230$, ${S}_{4}=0.080$, and ${S}_{5}=0.025$. In the learning of the $5$-qubit state, we achieve the fidelity ${f}\approx 0.9989$ by using $6.62\times 10^{7}$ success count. The result of state reconstruction is illustrated in Fig. 5 (b). Fig. 5 (c) shows the changes of infidelity during the learning process. The policy of RL agent induces exploration in the early time steps, and, as the ES approaches the optimal point, the agent focuses on exploitation to improve the quality of solution.
IV Discussion & Conclusion
The single-shot measurement scheme used in this work is proposed by the single-shot measurement learning (SSML) [25, 26]. Notably, the SSML addresses following problems of the standard approach: (i) The way to reconstruct a learned state is not known. (ii) The learned state can have negative eigenvalues [5]. (iii) The scaling of infidelity is limited to ${\cal O}(n^{-3/4})$ for $n$ copies of states without additional information about quantum state [73]. Our method also does not suffer from these problems. The trained HEA immediately provides a way to reconstruct the estimated state, and the reconstructed state always satisfies the conditions of quantum state as the state reconstruction is given by the quantum circuit. Also, the average infidelity nearly achieves the scaling of statistical limit [45]. While the SSML shows that the optimal scaling in the learning up to $6$-dimensional states, our results show the optimal scaling for higher dimensional states. Implementing the SSML on a quantum computer is difficult for high-dimensional states as it uses a parameterized special unitary gate and compiling the parameterized circuit typically incurs a high computational cost. In our method, we use the Hardware efficient ansatz of which structure can be compatible with quantum computers of low connectivity, e.g., one-dimensional array of qubits allowing neighborhood interaction.
A strategy of action repetition has been used in a reinforcement learning scheme for games [78]. In this context, the motivation of action repetition is to imitate human behaviors in playing games. On the other hand, we use the action repetition strategy to control the depth of decision process with the curriculum so that the agent can experience the problems of various difficulties. Reducing the depth and breadth of the decision process is a key issue in successfully applying reinforcement learning to large-scale problems [79]. (Note that, in our problem, the discretization of action space plays a role in reducing the breadth of decision process.) The efficacy of curriculum learning has been shown in meta-learning schemes for gradient descents [39] and the quantum optimization problems [36]. In the setting of our problem, the curriculum of the previous works can be realized by gradually increasing the maximum time steps of evolution strategy from a small value to a large value. However, we do not apply this curriculum to our method as the control of maximum time step directly affects the estimation accuracy of observation-value function. We estimate the value function based on the empirical cumulative reward, and we require the learning quantum state to be completed within a maximum time step to accurately compute the cumulative reward. If the maximum time step is set to a small value, the learning process will not be completed, and this hinders the calculation of cumulative reward. The inaccurate estimation of value function can bias the agent to only solve easy problem instances, the quantum states which can be learned in a short time step.
For the task of learning quantum states, we propose the meta-learning scheme based on the hybrid model of quantum and classical neural networks. We perform the learning Haar-random quantum states by training the Hardware efficient ansatz (HEA) of low connectivity with the evolution strategy. To improve the learning quantum states, we use the reinforcement learning (RL), and the RL agent learns the policy to dynamically adjust the hyperparameters of evolution strategy. The trained RL agent enables the infidelity of reconstructed states to approach the scaling of Heisenberg limit [45]. Our meta-learning model demonstrates the generalizability by showing that the RL agent can be applicable to learning quantum states of higher dimension than the dimension of quantum states used for training. The generalization is possible since the structure of RL agent is independent of the dimension of quantum system. To obtain these results, we introduce the action repetition strategy (ARS) which can facilitate the training of RL agent. We expect that our meta-learning scheme including the ARS can be leveraged to improve known classical and quantum algorithms for learning quantum states, quantum control [80], quantum machine learning [81], and quantum optimization [34, 82]. Furthermore, hyperparameter optimization and improving blackbox optimization algorithms through meta-learning have broad applications across disciplines [36, 83, 84, 85, 86].
Acknowledgements.
Authors thank Minyoung Lee and Taehee Lee for their support. J.J. thanks Sang Min Lee and Jeongho Bang for discussions.
Figure A1: The results of learning a thousand of random pure states obtained without the RL agent. The target success count is set to ${\cal C}_{\text{target}}=10^{4}$. We express the average of total success count and infidelity on a logarithmic scale. The total time steps are represented by ratio $t_{H}/t_{\text{max}}$, where $t_{\text{max}}=10^{4}$, $10^{5}$, and $10^{6}$ for $1$-, $2$-, and $3$-qubit state, respectively. The shaded cells are where the evolution strategy fails to learn quantum states within the maximum time step. The columns (a)–(c) are the results of the learning $1$-, $2$-, and $3$-qubit states with the $16$ actions, respectively. The column (d) is the result of the learning $2$-qubit states with the $169$ actions. The baseline values and corresponding action $a_{\text{base}}$ are represented by the red fonts.
Appendix A Markov decision process
For finite time steps $t=1,2,\ldots,T$, a markov decision process (MDP) is represented by a tuple $\left({\cal S},{\cal A},{\cal T},{\cal R}\right)$, where ${\cal S}$ is the space of state $s_{t}$, ${\cal A}$ is the space of action $a_{t}$, ${\cal T}$ is the space of transition probability $p(s_{t+1}|s_{t},a_{t})$ from a state $s_{t}$ to $s_{t+1}$ by an action $a_{t}$, and ${\cal R}$ is the space of reward $r_{t}$. In the description of MDP, a trajectory of decision process $\tau$, e.g., $s_{1},a_{1},r_{1},\ldots,s_{T-1},a_{T-1},s_{T},r_{T}$, is deemed to satisfy Markov condition, which assumes that a current state and reward only depend on a previous state and action. Specifically, the assumption implies that a probability distribution which governs the trajectory $\tau$ can be written as
$p(\tau)$ $=$ $p(s_{1},a_{1},r_{1},s_{2},a_{2},r_{2}\ldots,s_{T-1},a_{T-1},s_{T},r_{T})$ (A1)
$=$ $p(s_{1})\prod_{t=1}^{T-1}p(s_{t+1},r_{t+1}|s_{t},a_{t})p(a_{t}|s_{t}),$
where the probability $p(a_{t}|s_{t})$ is called policy. The purpose of agent is to learn the optimal policy so as to maximize the expected cumulative reward $G_{t}$ at each time step $t$, where
$G_{t}:={\mathbf{E}}_{\tau\sim p(\tau)}\left[\sum_{t}^{T}r_{t}\right].$ (A2)
Finally, the RL problem that the agent solves is an optimization to find a policy $\tilde{\pi}$ such that
$\tilde{\pi}=\mathop{\mathrm{argmax}}_{\{p(a_{t}|s_{t})\}}G_{0}.$ (A3)
Appendix B Elements of POMDP
Our RL problem can be formulated with a POMDP defined by a tuple $\left({\cal S},{\cal A},{\cal T},{\cal R},\Omega,{\cal O}\right)$:
  1. 1.
    ${\cal S}$ is the Hilbert space of $N$-qubit states. The state $s_{t}$ is a quantum state $\hat{U}(\vec{\theta}_{t})\left|\psi\right>$ transformed by the HEA in Fig. 1 (b).
  2. 2.
    ${\cal A}$ is the space of action $a_{t}=(\sigma_{t},\eta_{t})$, where $\sigma$ and $\eta$ are hyperprameters of evolution strategy which determine sampling range and learning rate, respectively [43].
  3. 3.
    ${\cal T}$ is the space of transition probability defined by $p(s_{t+1}|s_{t},a_{t}):{\cal S}\times{\cal A}\times{\cal S}\rightarrow[0,1]$.
  4. 4.
    ${\cal R}$ is the space of reward $r_{t}$. The reward $r_{t}=0$ if the learning quantum states is finished at time step $t$. Otherwise, $r_{t}=-1$.
  5. 5.
    $\Omega$ is the space of observation $o_{t}$. To observe the state, we use a measurement of binary outcome; success and fail. We perform the measurement until the fail outcome appears. The observation is the number of consecutive success outcomes before a fail outcome appears.
  6. 6.
    ${\cal O}$ is the space of the transition probability of observation defined by $p(o_{t}|s_{t}):{\cal O}\times{\cal S}\rightarrow[0,1]$.
Appendix C Actor-Critic algorithm
Actor-Critic algorithm proceeds with iteration of generating training datasets, value evaluation, and policy improvement [87, 88]. We generate the rollout datasets $\{(o_{t},a_{t},r_{t},o_{t+1})\}_{t=1}^{t_{H}-1}$ by using a tree search guided by the policy and store them in the data buffer. We randomly sample a datasets $\cal B$ from the data buffer [89], and, based on the sampled datasets $\cal B$, the critic network evaluates the observation-value function $Q_{\phi_{1}}$. We take the empirical cumulative reward (9) as an estimator of the observation-value function:
$\tilde{Q}(o_{t})=\frac{{\cal R}_{t}}{{t}_{\text{max}}},$ (A4)
where ${t}_{\text{max}}$ is the maximum time step of ES which is set to $3\times 10^{3}$, $10^{4}$, and $2\times 10^{4}$ for $1$-, $2$-, and $3$-qubit, respectively. We introduce ${t}_{\text{max}}$ in the denominator to normalize the empirical cumulative reward. The loss function of critic network is given by the difference between the output of critic network and the estimator of observation-value function, i.e.,
$loss_{\text{critic}}(\phi_{1}):=\frac{1}{2}\left(Q_{\phi_{1}}(o_{t})-\tilde{Q}(o_{t})\right)^{2}$ (A5)
We update the parameters of critic network $\phi_{1}$ by using a gradient descent method. Based on the critic network, we improve the policy $\pi_{\phi_{2}}$ by training the actor network. The loss function of actor network is given by
$loss_{\text{actor}}(\phi_{2}):={A(o_{t})}\log\pi_{\phi_{2}}(a_{t}|o_{t}),$ (A6)
where $A(o_{t}):=Q_{\phi_{1}}(o_{t})-\tilde{Q}(o_{t})$ is called advantage. The parameters of actor network $\phi_{2}$ is updated by a gradient ascent. In the typical Advantage Actor Critic (A2222C) method [90], the value function and advantage are estimated by using the time difference (TD) error. In our case, we estimate these quantities based on the empirical value function obtained by a policy-guided tree search, similar to RL using Monte-Carlo tree search [79, 91]. We use the ADAM optimizer for the gradient methods [92]. The learning rate of ADAM optimizer is set to $10^{-4}$ for the learning $1$-qubit states, and $3\times 10^{-5}$ for the learning $2$- and $3$-qubit states.
Appendix D Learning quantum states without RL agent
In Fig. A1, we investigate the performance of learning a thousand random pure states without using the RL agent. We choose the performance given by an action $a_{\text{base}}$ which gives the lowest total success count as the baseline. For $1$-qubit, $a_{\text{base}}=(\sigma=0.1,\eta=0.01)$. In case of $2$-qubit, $a_{\text{base}}=(\sigma=0.1,\eta=0.001)$ for the $16$ actions, and $a_{\text{base}}=(\sigma=10^{-1},\eta=10^{-2.75})$ for the $169$ actions. For $3$-qubit, $a_{\text{base}}=(\sigma=0.033,\eta=0.0001)$. The respective baseline values are represented by the red fonts in Fig. A1, and the dashed lines in Fig. 2. The shaded cells are where the evolution strategy fails to learn quantum states within the maximum time step $t_{\text{max}}$. These results imply that, without the RL agent, finding an action which can successfully complete the learning quantum states requires much amount of computational resources. For example, without the trained RL agent, learning a thousand $3$-qubit states requires more than a million time steps for most of the actions considered.